Xolisile and Xolani are twins. To reward them for passing matric well, their parents gave them R20 000,00 each. Each one of them decided to invest three quarters of their money for a period of three years. Xolisile invested her money at a simple interest rate of 6,5% per annum. Xolani invested her money at an interest rate of 6,5% per annum, compounded annually.  What is the difference between the values of the two investments at the end of the period?


1
Expert's answer
2021-06-03T12:46:11-0400

Xolasile.

P = (3/4)* R 20000

= R 15,000.

Rate = 6.5%

Time = 3 years

Interest = p*r*t

= 15000*(6.5/100)*3

= 15000*0.065*3

= R 2,925

Total Amount = R 2,925 + R 15,000

= R 17,925


Xolani

Principal = 0.75*20,000

= R 15,000

Rate = 6.5 %

Time=3 years

A = p(1+r/100)t

15000(1+0.065)3

15000(1.065) 3

15000*1.208

= R 18,119.24

Xolani has a higher earning at the end of 3 years with (R 18,119.24. - R 17,925)

= R 194.24